# Interpreting token compositionality in LLMs: A robustness analysis

### v1

@@ 5 Information Gain & Token Mutual Information
$IG_{R_{l_{p}}(X),R_{l_{q}}(X)}<IG_{R_{l_{p}}(X),R_{l_{r}}(X)},\quad\forall p<q<r$ (5)
where $p$, $q$ and $r$ are layer indices, i.e., later layers have more information about the inputs than earlier ones. This can be explained in that optimising Equation 4 can be achieved by retaining at each $R_{l_{p}}(X)$ only the necessary information to maximise $\sum_{i,j}{IG_{R_{l_{q}}(X_{i}),MHA(R_{l_{p}}(X_{j}))}}$, where $MHA(R_{l_{p}}(X_{j}))$ is the multi-head attention weighted representation. Such an objective implies minimising the mutual information $I(R_{l_{p}}(X_{i}),R_{l_{p}}(X_{j}))$, i.e., reducing redundancy across tokens from the same layer. Therefore, token dependencies will tend to be modelled by aggregation paths spanning multiple layers, with more layers allowing for more complex and longer paths. This is in line with the findings of Mechanistic Interpretability studies Elhage et al. (2021); Conmy et al. (2023). Equation 5 also implies that the earlier an aggregation is done, the larger the impact it will have on $IG_{Y,X}$, which explains the empirical results.
@@ 6 Related work
General compositionality
Compositionality is a core concept across multiple disciplines that refers to understanding complex systems through the meanings of their simpler components. It appears in philosophy, cognitive science, linguistics, and machine learning Tull et al. (2024). In machine learning (ML) and artificial intelligence (AI), compositionality enables their models to generalise to, reason about, and interpret complex meanings from basic, known components. In cognitive science, the Language of Thought Hypothesis (LOTH) Fodor (1975) hypothesises that mental processes are compositional, constructing complex thoughts from basic elements Rescorla (2024). Recent ML studies explore how transformers and neural networks build their inferences, dissecting potential compositionality. Dai et al. (2022) show that feedforward layers in transformers act as key-value memories, where the models combine information to create complex hashing-like predictions. MEMIT Meng et al. (2023) and PMET Li et al. (2024) reveal how controlled inferences can be built by manipulating models’ parts. These studies suggest transformers may employ compositional processing, systematically building complex representations from simpler components.
Linguistic compositionality.
In linguistics, compositionality is often linked to the principle of compositionality, which states that the meaning of a complex expression is derived from "the meanings of its constituent parts, the grammar used to combine them, and the syntactic structure as a whole"Montague and Thomason (1975). Yu and Ettinger (2020) found that transformers mainly encode individual word content rather than true phrase-level meaning, with model performance often tied to word overlap rather than compositional understanding. Carvalho et al. (2022) probed transformer models’ representation of adjectival modifier phenomena in adjective-noun phrases. Their tests revealed that models partially capture meaning intersective composition, which does not generalise across other adjective types. DecompX, a method introduced in Modarressi et al. (2023), aims to explain transformer models by decomposing token representations and tracing their aggregation through all layers. It provides a more faithful explanation of model predictions by linking individual token dynamics to final decisions without the need for vector aggregation across layers, which offers compositional behaviour analysis on a token-prediction basis in transformer models. Haslett and Cai (2024), examine how LLMs segment words into multiple tokens and the extent of semantic information encoded within these tokens. Their findings show that models from Multilingual BERT and GPT-2 to GPT-4o often fail to segment words meaningfully, particularly in non-Latin script languages, where morphemes are not well-represented.
This work complements and expands the aforementioned studies’ findings to characterise the previously observed phenomena and better understand the disconnection between model representations and compositional semantics reported in all of them.
@@ 7 Conclusion
This work presents a systematic analysis of the robustness of Transformer-based LLMs to perturbations grounded in compositionality. Motivated by prior studies highlighting a significant and unexpected gap between linguistic compositionality and LLM representations, we aimed to characterise the impact of compositional aggregation at each inference step and provide an information-theoretical explanation for the observed results. Our findings indicate a clear pattern of information structuring across contemporary LLMs, where token dependencies tend to be modelled by aggregation paths spanning multiple layers, and the capability to learn complex token structures is gained at the cost of higher sensitivity to perturbations at the inputs and earlier layers. Furthermore, based on the relation between the information gain from input representation to predicted token and the mutual information between token representations in the model, we postulate that compositional semantic representations do not exist in any particular stage of a standard Transformer model for a general input. These findings point to future work on compositional-aware models towards the use of specialised architectures and/or training objectives to elicit such representations.
@@
A number of limitations need to be acknowledged regarding our paper. First, the dataset used, WordNet, may not fully represent the diversity of language usage across all relevant domains. Secondly, the employed transformer models are not exempted from possible biases or limitations in their training data or modelling. Third, our findings rely on the Benepar parsing model, which is liable to inaccuracies as far as linguistic analysis is concerned. Thus, the accuracy of our conclusions is limited to the parsing model performance. Furthermore, the applicability of our results to other languages has not been tested. In future research, these shortcomings should be addressed, and it might also be possible to combine languages and parsing models to further validate and extend our findings.
@@ Appendix A Compositionality and Localisation
The concept of linguistic compositionality has evolved from its origins in Frege’s work Frege (1892), which started conceptualising the notion that the meaning of a complex expression is determined by its constituent parts and their syntactic arrangement. This principle was formalized by Montague Montague (1970b, a), who applied mathematical rigour to natural language semantics, thereby reinforcing the compositional approach within formal semantics. Linguistic phenomena such as idioms, context-dependence, and metaphor, which seemed to violate compositionality, prompted debates on its universality Katz and Postal (1963); Jackendoff (1997), with theoretical accounts evolving to integrate these phenomena, leading to a more nuanced understanding that balances strict compositional rules with allowances for non-compositional elements Partee (1984).
@@ B.1 Downstream Task Definitions
Inverse Definition Modelling (IDM): The Inverse Definition Modelling (IDM) task involves predicting a term $T$ based on a given natural language definition $D$. Let $D=\{d_{1},d_{2},\dots,d_{n}\}$ represent the sequence of tokens constituting the definition. The goal is to generate the corresponding term $T$, where:
@@
Synonym Prediction (SP): The Synonym Prediction (SP) task requires the model to generate a synonym $S$ for a given word $W$. Let $W\in\mathcal{V}$ represent the input word. The task is to predict a synonym $S$, such that:
@@
Hypernym Prediction (HP): The Hypernym Prediction (HP) task involves predicting a more general term, or hypernym, $H$ for a given word $W$. Let $W\in\mathcal{V}$ represent the input word. The objective is to predict a hypernym $H$, such that:
@@ B.3 Model Specifications and Fine-tuning Parameters
The results in Table 3 provide a comparative overview of various Transformer models used in this study. We used GPT-2 models (released under the Modified MIT License), Gemma-2B (released under the Gemma Terms of Use), and Llama3-8B (released under the Meta Llama 3 Community License). All models were used for research purposes, specifically for language modeling and text generation in English, aligning with their intended usage. The models differ in their number of parameters, layers, heads, and feedforward (FF) dimensions. The number of parameters ranges from 85M for GPT2-small to 7.8B for LLama3-8B. The activation functions and FF dimensions also highlight variations in the internal processing architecture, influencing the models’ performance across different tasks. In addition to these architectural differences, the models were fine-tuned using a consistent set of hyperparameters. The fine-tuning process spanned over 3 training epochs with a batch size of 16. The learning rate was set to 5e-5, while a weight decay of 0.01 was applied to prevent overfitting. Training logs were generated every 200 steps, with model checkpoints saved every 1000 steps, but limited to retaining only one checkpoint to manage storage efficiently. The evaluation strategy during fine-tuning was set to evaluate at the end of each epoch, and similarly, the model was saved at the end of each epoch as well.
Appendix C Detailed Performance Evaluation and Results
C.1 Baseline Performance
@@
C.2 Comprehensive CAP Results for All Models and Tasks
Table 5 and Table 6 present the reduction in accuracy when applying word-level and phrasal CAP, respectively, to several models—GPT2-small, GPT2-medium, GPT2-large, Gemma-2B, and Llama3-8B—across three tasks: Inverse Dictionary Modelling (IDM), Synonym Prediction (SP), and Hypernym Prediction (HP). The results of phrasal-level CAP for Gemma-2B and Llama3-8B are not reported due to the severe degradation in model performance under these conditions, rendering the outputs effectively unusable.
@@ C.2 Comprehensive CAP Results for All Models and Tasks
Table 5: Performance drop (in percentage points) for GPT2-small, GPT2-medium, GPT2-large, Gemma-2B, and Llama3-8B models after applying word-level CAP across three tasks: Inverse Dictionary Modelling (IDM), Synonym Prediction (SP), and Hypernym Prediction (HP). Results are reported for different layer positions (1%, 25%, 75%, and 100%) in both Original and Fine-tuned settings, using three CAP protocols: Max, Mean, and Sum.

### v2

@@ 5 Information Gain & Token Mutual Information
$IG_{R_{l_{p}}(X),R_{l_{q}}(X)}<IG_{R_{l_{p}}(X),R_{l_{r}}(X)},\quad\forall p<q<r,$ (6)
where $p$, $q$ and $r$ are layer indices, i.e., subsequent layers have more information about the inputs than previous ones. This can be explained in that optimising Equation 5 can be achieved by retaining at each $R_{l_{p}}(X)$ only the necessary information to maximise $\sum_{i,j}{IG_{R_{l_{q}}(X_{i}),MHA(R_{l_{p}}(X_{j}))}}$, where $MHA(R_{l_{p}}(X_{j}))$ is the multi-head attention weighted representation. Such an objective implies minimising the mutual information $I(R_{l_{p}}(X_{i}),R_{l_{p}}(X_{j}))$, i.e., reducing redundancy across tokens from the same layer. Therefore, token dependencies will tend to be modelled by aggregation paths spanning multiple layers, with more layers allowing for more complex and longer paths. This is in line with the findings of Mechanistic Interpretability studies Elhage et al. (2021); Conmy et al. (2023). Equation 6 also implies that the earlier an aggregation is done, the larger the impact it will have on $IG_{Y,X}$, which explains the empirical results.
@@ 6 Related work
General compositionality. Compositionality, a fundamental concept across disciplines, concerns understanding complex systems through their simpler components’ meanings. It spans philosophy, cognitive science, linguistics, and machine learning Tull et al. (2024). In ML and AI, compositionality enables models to generalise, reason about, and interpret complex meanings from basic components. In cognitive science, the Language of Thought Hypothesis (LOTH) Fodor (1975) proposes that mental processes are compositional Rescorla (2024). Recent ML studies explore how transformers and neural networks build inferences. The logit lens Nostalgebraist (2020) demonstrated that transformers build predictions progressively where early layers make initial guesses and deeper layers refine guesses with broader context. Dai et al. (2022) show feed-forward layers act as key-value memories, combining information for complex predictions. MEMIT Meng et al. (2023) and PMET Li et al. (2025) show how controlled inferences can be built by manipulating models’ components. These studies suggest transformers may employ compositional processing, systematically building complex representations from simpler components.
Linguistic compositionality. In linguistics, compositionality is often linked to the principle of compositionality, which states that the meaning of a complex expression is derived from "the meanings of its constituent parts, the grammar used to combine them, and the syntactic structure as a whole" Montague and Thomason (1975). Yu and Ettinger (2020) found that transformers mainly encode individual word content rather than true phrase-level meaning, with model performance often tied to word overlap rather than compositional understanding. Carvalho et al. (2025) probed transformer models’ representation of adjectival modifier phenomena in adjective-noun phrases. Their tests revealed that models partially capture meaning intersective composition, which does not generalise across other adjective types. DecompX Modarressi et al. (2023) traces individual token representations through transformer layers, enabling analysis of compositional behaviour at the level of specific token predictions without requiring cross-layer vector aggregation. Haslett (2024) demonstrate that models from Multilingual BERT through GPT-4 frequently fail to segment words meaningfully, especially in non-Latin scripts where morpheme boundaries are poorly captured.
This work complements and extends these findings to characterise the fundamental disconnect between transformer representations and compositional semantics observed across aforementioned studies.
@@ 7 Conclusion
This work systematically analyses the robustness of transformer-based LLMs to compositional perturbations. Motivated by studies highlighting an unexpected gap between linguistic compositionality and LLM representations, we characterised the impact of compositional aggregation at each inference step and provided an information-theoretical explanation. Our findings indicate a pattern where token dependencies are modelled by aggregation paths spanning multiple layers, and complex token structure learning comes at the cost of higher sensitivity to perturbations at inputs and earlier layers. Based on the relation between information gain from input to predicted token and mutual information between token representations, we postulate that compositional semantic representations cannot be isolated to any particular (intermediate) stage of a standard transformer model. These insights suggest that future compositional-aware models should explore specialised architectures or training objectives. Natural extensions include analysing encoder-based and encoder-decoder transformers and investigating final token representations to further understand internal compositional mechanisms.
@@
Several limitations are acknowledged in our paper. First, the WordNet dataset may not fully represent language diversity across all domains. Second, the employed transformer models are decoder-based only and could be subject to biases from their training data. Third, our findings depend on the Benepar parsing model, which may introduce inaccuracies in linguistic analysis. Additionally, the applicability of our results to other languages has not been tested. Future research should address these issues and consider combining languages and parsing models for further validation.
@@ Appendix A Compositionality and Localisation
The concept of linguistic compositionality has evolved from its origins in Frege’s work Frege (1892), which started conceptualising the notion that the meaning of a complex expression is determined by its constituent parts and their syntactic arrangement. This principle was formalised by Montague Montague (1970b, a), who applied mathematical rigour to natural language semantics, thereby reinforcing the compositional approach within formal semantics. Linguistic phenomena such as idioms, context-dependence, and metaphor, which seemed to violate compositionality, prompted debates on its universality Katz and Postal (1963); Jackendoff (1997), with theoretical accounts evolving to integrate these phenomena, leading to a more nuanced understanding that balances strict compositional rules with allowances for non-compositional elements Partee (1984).
@@ B.1 Downstream Task Definitions
Inverse Definition Modelling (IDM): The IDM task involves predicting a term $T$ based on a given natural language definition $D$. Let $D=\{d_{1},d_{2},\dots,d_{n}\}$ represent the sequence of tokens constituting the definition. The goal is to generate the corresponding term $T$, where:
@@
Synonym Prediction (SP): The SP task requires the model to generate a synonym $S$ for a given word $W$. Let $W\in\mathcal{V}$ represent the input word. The task is to predict a synonym $S$, such that:
@@
Hypernym Prediction (HP): The HP task involves predicting a more general term, or hypernym, $H$ for a given word $W$. Let $W\in\mathcal{V}$ represent the input word. The objective is to predict a hypernym $H$, such that:
@@ B.2 Dataset Descriptions and Preprocessing
LLama3-8B (Instruct) 7.8B 32 4096 32 SiLU 14336
Qwen2.5-0.5B 391M 24 896 14 SiLU 4864
Qwen2.5-1.5B 1.4B 28 1536 12 SiLU 8960
Qwen2.5-3B 3.0B 36 2048 16 SiLU 11008
@@
Llama3-8B (Instruct) 1.61% 8.41% 12.09% 19.36% 10.97% 24.73%
Llama3-3B 20.51% 8.26% 12.19% 26.42% 13.43% 31.1%
Qwen-0.5B 8.21% 6.10% 12.03% 18.83% 10.94% 28.03%
Qwen-1.5B 12.35% 7.61% 14.64% 30.01% 13.70% 31.31%
Qwen-3B 13.35% 7.53% 14.40% 31.80% 13.66% 31.95%
@@ B.3 Model Specifications and Fine-tuning Parameters
The results in Table 3 provide a comparative overview of various Transformer models used in this study. We used GPT-2 models (released under the Modified MIT License), Gemma-2B (released under the Gemma Terms of Use), Llama3 models (released under the Meta Llama 3 Community License), and Qwen models (released under Apache License 2.0). All models were used for research purposes, specifically for language modelling and text generation in English, while Qwen models provide additional multilingual support, aligning with their intended usage. The models differ in their number of parameters, layers, heads, and feed-forward (FF) dimensions. The number of parameters ranges from 85M for GPT2-small to 7.8B for LLama3-8B. The activation functions and FF dimensions also highlight variations in the internal processing architecture, influencing the models’ performance across different tasks. In addition to these architectural differences, the models were fine-tuned using a consistent set of hyperparameters. The fine-tuning process spanned over three training epochs with a batch size of 16. The learning rate was set to 5e-5, while a weight decay of 0.01 was applied to prevent overfitting. Training logs were generated every 200 steps, with model checkpoints saved every 1000 steps, but limited to retaining only one checkpoint to manage storage efficiently. The evaluation strategy during fine-tuning was set to evaluate at the end of each epoch, and similarly, the model was saved at the end of each epoch as well.
Appendix C Token Reduction Analysis
Table 5 presents an analysis of activation reduction percentages across different LLMs, particularly for the token-to-words case. In this context, the mean represents the average reduction percentages across samples, while the standard deviation indicates the variability of these reductions. The purpose is to assess whether token reduction across models would highly influence the results of CAP.
Token reduction is a factor but not the sole determinant of performance degradation.
The results presented in Tables 7,  8, and 9 indicate that while token reduction percentage influences performance degradation, it is not the sole determining factor. Several key observations support this conclusion, which is discussed below.
Model Task Mean ± Std
GPT2 (S) IDM $3\pm 5$
SP $27\pm 9$
HP $27\pm 10$
GPT2 (M) IDM $3\pm 5$
SP $28\pm 10$
HP $26\pm 11$
GPT2 (L) IDM $3\pm 5$
SP $27\pm 9$
HP $26\pm 11$
Gemma-2B IDM $9\pm 4$
SP $19\pm 9$
HP $30\pm 9$
Llama3-3B IDM $10\pm 5$
SP $23\pm 6$
HP $28\pm 6$
Llama3-8B IDM $10\pm 5$
SP $21\pm 7$
HP $28\pm 9$
Llama3-8B instruct IDM $13\pm 4$
SP $30\pm 6$
HP $28\pm 8$
Qwen 0.5B IDM $3\pm 5$
SP $9\pm 11$
HP $20\pm 10$
Qwen 1.5B IDM $3\pm 5$
SP $12\pm 10$
HP $19\pm 10$
Qwen 3B IDM $3\pm 5$
SP $12\pm 10$
HP $19\pm 10$
Table 5: Reduction percentages
First, we observe that higher token reduction does not always lead to a greater performance drop. For instance, models such as Gemma-2B and Llama3-8B exhibit high token reduction percentages (Table 5), yet their performance degradation varies significantly across tasks and layer positions. Also, despite lower token reduction percentages, the models Qwen 0.5B and GPT2-small still show substantial accuracy drops, particularly in early layers in the SP and HP tasks. Second, model size and depth influence degradation, as evident in the larger models (e.g., Llama3-8B, Gemma-2B) exhibiting greater fragility to CAP interventions, particularly in early layers (1% and 25%). Third, as discussed in the paper, layer-specific variability suggests hierarchical processing differences. Early-layer CAP interventions cause severe accuracy drops in large models but have a less pronounced effect in smaller models, suggesting that deeper architectures defer compositional integration to later layers. Further, fine-tuning reduces degradation in later layers (75% and 100%), implying that learned representations in deeper layers mitigate the effects of early perturbations. Finally, architectural differences influence sensitivity. We observe that higher MLP dimensions (e.g., Llama3-8B: 14,336 vs. GPT2-small: 3,072) correlate with greater vulnerability to CAP perturbations, likely due to increased parameter redundancy and disruption of the key-value recall mechanism in MLPs Meng et al. (2022).
While the token reduction percentage contributes to performance degradation, it is insufficient to fully explain the observed variations. Task nature, model size, layer depth, activation functions, and MLP dimensions collectively influence the robustness of CAP interventions. Larger, deeper models demonstrate greater sensitivity to early perturbations, while fine-tuning helps recover performance in later layers. These findings suggest that effective compositional representations in LLMs are distributed rather than localised, requiring specialised architectures or training objectives to improve robustness.
Appendix D Evaluating Parsing Accuracy and Addressing the Impact of Benepar Parser Errors
A key potential bias in our results comes from the reliance on the constituency parser for token-to-phrase experiments. Inaccuracies in parsing may distort the results of CAP. To address this, we report the chosen parser’s accuracy by testing it on the Stanford Sentiment Treebank (SST) dataset, a dataset that offers golden labels for parsing. We aim to alleviate concerns about the parser’s impact on our findings by showcasing its accuracy on the SST dataset. The parser evaluation was conducted as follows:
Dataset.
A subset of 1,000 randomly sampled sentences from the test split of the SST dataset was used for the analysis. The Stanford Sentiment Treebank (SST) provides annotated constituency labels, which serve as the golden labels for comparison with parser outputs. While WordNet definitions offer rich semantic information, they lack annotated golden constituency labels, making direct parser validation infeasible. The use of SST’s annotations enables reliable parser evaluation and indirectly supports the validation of the parsing correctness for WordNet definitions, provided they follow standard syntactic structures.
Parser.
The Benepar parser was employed for parsing sentences due to its strong performance in constituency parsing tasks. Benepar is widely recognised for its robustness and ability to handle diverse syntactic structures. For this evaluation, the constituency structures generated by Benepar were directly compared against SST’s golden annotations to assess its parsing accuracy.
Evaluation metrics.
The parser’s performance was evaluated using the following metrics:
  • Precision: Proportion of correctly predicted constituents out of all predicted constituents.
  • Recall: Proportion of correctly predicted constituents out of all ground truth constituents.
  • F1-Score: Harmonic mean of precision and recall, providing an overall performance measure.
  • Accuracy: Percentage of sentences where the predicted constituency structure fully matches the ground truth.
Results robustness.
To ensure robustness and consistency, the evaluation was repeated across five different random seeds. This allowed for an assessment of variability in performance across multiple subsets of the dataset. Additionally, constituents were evaluated at hierarchical levels—such as root level, phrase level, and token level—to analyse parsing performance across varying syntactic granularities.
Results.
The evaluation yielded the following averaged metrics across five seeds for the default level of parsing (Level 1, the immediate children of the root node):
Metric Mean ± Std
Precision 0.956 ± 0.001
Recall 0.956 ± 0.001
F1-Score 0.956 ± 0.001
Accuracy 0.956 ± 0.001
Table 6: Aggregated evaluation metrics for Level 1 constituents using the Benepar parser, averaged across five seeds.
Interpretation.
The results demonstrate consistently high parsing accuracy across all evaluation metrics, with minimal variability (as indicated by the low standard deviation). These findings validate the Benepar parser’s reliability for parsing Level 1 constituents, which form the backbone of sentence structure. Consequently, the parser’s impact on CAP results is minimal, ensuring robustness and validity of our conclusions.
Appendix E Detailed Performance Evaluation and Results
E.1 Baseline Performance
@@
E.2 Comprehensive CAP Results for All Models and Tasks
Tables 7, 8, 9, 10, 11, and 12 present the reduction in accuracy when applying word-level and phrasal CAP, respectively, across models and the three tasks: Inverse Dictionary Modelling (IDM), Synonym Prediction (SP), and Hypernym Prediction (HP). The results of phrasal-level CAP for Gemma-2B and Llama3-8B are not reported due to the severe degradation in model performance under these conditions, rendering the outputs effectively unusable.
@@ E.2 Comprehensive CAP Results for All Models and Tasks
Llama3-3B 1% 28.79% 26.36% 25.96% 25.54% 22.71% 22.74%
25% 31.73% 8.08% 6.99% 13.44% 5.84% 5.8%
75% 12.27% 5.84% 5.22% 8.54% 5.03% 5.15%
100% 3.62% 1.99% 1.99% 2.37% 1.82% 1.85%
Llama3-8B Instruct 1% 41.75% 48.99% 45.99% 22.01% 19.97% 20.29%
25% 44.43% 25.22% 25.12% 17.62% 13.51% 15.22%
75% 26.17% 24.42% 24.98% 7.9% 5.34% 5.19%
100% 21.8% 12.04% 12.04% 1.51% 1.29% 1.33%
Qwen2.5-0.5B 1% 10.12% 8.2% 8.23% 7.85% 6.39% 6.00%
25% 5.19% 4.21% 4.45% 4.35% 3.29% 3.49%
75% 3.56% 2.82% 3.15% 2.39% 2.24% 2.15%
100% 0.98% 0.98% 0.98% 0.23% 0.28% 0.33%
Qwen2.5-1.5B 1% 14.56% 11.04% 10.22% 9.47% 7.36% 7.48%
25% 13.29% 4.45% 5.34% 6.83% 3.86% 4.00%
75% 7.03% 2.68% 2.84% 4.21% 2.74% 2.79%
100% 0.7% 0.4% 0.4% 0.65% 0.23% 0.23%
Qwen2.5-3B 1% 12.63% 12.27% 11.44% 7.85% 6.71% 6.48%
25% 18.61% 8.59% 9.11% 10.66% 4.75% 5.82%
75% 7.23% 4.00% 3.79% 3.65% 2.83% 2.8%
100% 0.39% 0.4% 0.4% 0.31% 0.17% 0.2%
Table 7: Performance drop (in percentage points) for GPT2 (small, medium, large), Gemma-2B, Llama3 (3B, 8B, 3B Instruct), and Qwen2.5 (0.5B, 1.5B, 3B) models after applying word-level CAP for the Inverse Dictionary Modelling (IDM) task. Results are reported for different layer positions (1%, 25%, 75%, and 100%) in both Original and Fine-tuned settings, using three CAP protocols: Max, Mean, and Sum.
Model Layer Position Original Fine-tuned
Max Mean Sum Max Mean Sum
@@
Llama3-3B 1% 100% 100% 100% 100% 100% 100%
25% 85.81% 86.2% 85.16% 88.47% 84.54% 85.48%
75% 40.18% 39.3% 38.91% 14.77% 16.48% 15.64%
100% 5.77% 6.16% 6.16% 5.8% 6.12% 6.12%
Llama3-8B (Instruct) 1% 85.46% 85.71% 83.88% 42.9% 32.33% 31.09%
25% 54.34% 55.48% 58.83% 27.52% 18.13% 18.27%
75% 44.92% 31.51% 25.77% 19.23% 12.5% 12.23%
100% 6.81% 7.3% 7.3% 5.77% 4.81% 4.95%
Qwen2.5-0.5B 1% 81.77% 88.89% 79.17% 64.24% 58.36% 53.3%
25% 90.8% 91.15% 86.11% 54.51% 54.38% 37.22%
75% 63.72% 66.32% 39.06% 48.87% 48.57% 24.29%
100% 8.51% 10.07% 8.51% 3.67% 3.8% 3.8%
Qwen2.5-1.5B 1% 89.35% 84.52% 84.23% 64.55% 56.79% 56.03%
25% 90.58% 83.48% 83.19% 60.45% 55.5% 54.79%
75% 22.06% 22.21% 18.8% 10.88% 10.34% 10.02%
100% 6.82% 3.55% 3.55% 8.19% 7.87% 7.87%
Qwen2.5-3B 1% 81.39% 81.53% 73.58% 55.93% 49.35% 49.57%
25% 93.04% 89.91% 82.81% 72.41% 42.78% 38.47%
75% 77.84% 69.6% 49.43% 43.24% 22.13% 15.25%
100% 3.98% 3.13% 3.13% 1.4% 1.29% 1.29%
Table 8: Performance drop (in percentage points) for GPT2 (small, medium, large), Gemma-2B, Llama3 (3B, 8B, 3B Instruct), and Qwen2.5 (0.5B, 1.5B, 3B) models after applying word-level CAP for the Synonym Prediction (SP) task. Results are reported for different layer positions (1%, 25%, 75%, and 100%) in both Original and Fine-tuned settings, using three CAP protocols: Max, Mean, and Sum.
Model Layer Position Original Fine-tuned
Max Mean Sum Max Mean Sum
@@
Llama3-3B 1% 100% 99.95% 99.95% 99.93% 99.86% 99.82%
25% 88.04% 83.87% 84.34% 65.53% 63.92% 64.17%
75% 26.06% 24.47% 23.4% 11.06% 10.52% 10.79%
100% 4.34% 4.31% 4.31% 3.85% 4.08% 3.86%
Llama3-8B (Instruct) 1% 92.51% 92.45% 92.48% 95.86% 96.21% 96.03%
25% 68.41% 65.9% 67.24% 70.72% 69.99% 69.64%
75% 15.79% 14.78% 15.55% 20% 20.29% 19.53%
100% 0.57% 0.39% 0.39% 3.57% 3.5% 3.29%
Qwen2.5-0.5B 1% 93.76% 90.95% 85.27% 86.33% 80.55% 77.91%
25% 97.12% 97.51% 89.18% 74.83% 75.41% 75.77%
75% 76.74% 77.96% 55.39% 50.69% 49.71% 48.81%
100% 6.15% 5.56% 5.56% 2.48% 2.34% 2.34%
Qwen2.5-1.5B 1% 97.14% 90.5% 88.96% 88.52% 83.19% 77.21%
25% 98.12% 95.66% 94.04% 72.29% 68.18% 68.33%
75% 18.27% 18.72% 17.94% 8.94% 9.64% 9.51%
100% 7.13% 6.81% 6.81% 3.95% 3.8% 3.8%
Qwen2.5-3B 1% 83.26% 82.41% 68.8% 75.13% 72.56% 70.69%
25% 97.36% 96.32% 88.81% 92.69% 79.67% 79.63%
75% 86.56% 71.45% 45.47% 40.87% 30.95% 33.04%
100% 2.07% 1.89% 1.89% 0.45% 0.35% 0.41%
Table 9: Performance drop (in percentage points) for GPT2 (small, medium, large), Gemma-2B, Llama3 (3B, 8B, 3B Instruct), and Qwen2.5 (0.5B, 1.5B, 3B) models after applying word-level CAP for the Hypernym Prediction (HP) task. Results are reported for different layer positions (1%, 25%, 75%, and 100%) in both Original and Fine-tuned settings, using three CAP protocols: Max, Mean, and Sum.
@@
Qwen2.5-0.5B 1% 93.97% 91.19% 87.15% 90.94% 84.44% 78.85%
25% 84.64% 76.78% 78.00% 76.36% 66.24% 67.16%
75% 61.75% 57.95% 63.86% 48.86% 41.8% 46.25%
100% 32.29% 26.8% 19.5% 13.55% 10.17% 15.08%
Qwen2.5-1.5B 1% 98.24% 95.8% 95.82% 93.31% 87.33% 80.81%
25% 96.4% 84.72% 89.41% 79.52% 63.00% 65.53%
75% 69.68% 64.6% 60.33% 19.11% 14.72% 24.01%
100% 68.03% 60.04% 56.6% 12.01% 7.46% 12.72%
Qwen2.5-3B 1% 96.51% 94.37% 94.64% 90.11% 86.02% 80.57%
25% 96.82% 89.89% 92.39% 90.24% 76.55% 76.28%
75% 82.27% 74.71% 77.07% 47.45% 36.06% 39.95%
100% 62.26% 62.21% 58.12% 7.41% 5.52% 8.18%
Model Layer Position Original Fine-tuned
Max Mean Sum Max Mean Sum
@@
Qwen2.5-0.5B 1% 99.32% 95.88% 92.87% 81.67% 61.89% 57.95%
25% 98.65% 95.91% 96.45% 60.19% 58.75% 58.43%
75% 93.21% 84.66% 77.4% 56.29% 49.3% 44.94%
100% 68.78% 45.74% 43.92% 13.56% 7.47% 16.79%
Qwen2.5-1.5B 1% 98.1% 96.33% 94.43% 72.33% 58.5% 59.55%
25% 97.55% 96.2% 95.38% 63.79% 55.84% 68.93%
75% 75.72% 55.17% 48.41% 19.33% 14.48% 26.87%
100% 70.39% 38.68% 36.29% 18.73% 10.41% 20.97%
Qwen2.5-0.5B 1% 96.47% 95.52% 90.31% 74.05% 67.1% 56.57%
25% 99.32% 98.1% 94.29% 94.89% 56.93% 57.38%
75% 94.02% 89.46% 83.4% 86.43% 64.01% 43.39%
100% 47.00% 35.56% 31.32% 20.07% 15.19% 21.15%
Model Layer Position Original Fine-tuned
Max Mean Sum Max Mean Sum